# Systematic bias in LISA ringdown analysis due to waveform inaccuracy

### v1

@@ Abstract
[…]
@@ I Introduction
As a binary black hole (BH) system emits gravitational waves (GWs), the two black holes spiral closer together, ultimately merging in a coalescence event. Immediately after the merger, the remnant BH is in a highly perturbed state and continues to emit GWs in a ringdown phase, gradually settling into a stable configuration. The GW signal in this regime is well understood in terms of BH perturbation theory. In general, any perturbation of the spacetime metric of a rotating BH in General Relativity (GR), at linear order, is governed by the Teukolsky equation Teukolsky (1973). Solutions to the Teukolsky equation, given appropriate boundary conditions, are expressed as a linear superposition of quasi-normal modes (QNMs), i.e. exponentially damped sinusoids Vishveshwara (1970); Chandrasekhar and Detweiler (1975); Anninos et al. (1993); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). Specifically, QNMs are characterized by complex frequencies and amplitudes, with the imaginary part of the frequency determining the damping timescale.
@@ II.1 Time and frequency-domain models
where $M$ is the remnant BH mass in the detector frame, $d_{L}$ is the luminosity distance to the source, $\omega_{\ell mn}$ and $-1/\tau_{\ell mn}$ are the real and imaginary parts of the QNM frequency, and $(\iota,\varphi)$ are the angular coordinates of the direction of propagation as seen from the remnant BH, with $\prescript{}{-2}{Y_{\ell m}}$ the spin-weighted spherical harmonics. Note that for a given set of $(\ell mn)$, there are two mode contributions $\mathcal{A}_{\ell mn}$ and $\mathcal{A}^{\prime}_{\ell mn}$: this is because for any given QNM with frequency $\omega_{\ell mn}$ and damping time $\tau_{\ell mn}$, there exists a “mirror mode” also solving the equations of motion, which corresponds to the second term in Eq. (II.1Berti et al. (2009, 2006b). Finally, for this study, we set the starting time to $t_{0}=20M$ after the luminosity peak. For further details on the choice of the starting time, see App. B
@@
where * denotes complex conjugation. This property follows from equatorial symmetry Bucciotti et al. (2024a); Isi and Farr (2021), i.e. from neglecting orbital precession in the merger waveform.
@@
×$\times\sum_{\ell mn}\frac{A_{\ell mn}Y_{\ell m}^{+,\times}}{\tau_{\ell mn}}\left(L_{\ell mn}^{+}\pm L_{\ell mn}^{-}\right)\,,$ (4)
@@
Given two times series $u(t)$ and $v(t)$, we define the matched filter inner product as
@@ II.2 Validity regime of the QNM model
To assess the number of modes needed to ensure waveform accuracy, it is important to account for the validity regime of the frequency-domain QNM model given by Eq. (II.1). The time-domain damped-sinusoid QNM model given by Eq. (II.1) is only accurate at intermediate times after the merger, because early times are dominated by the prompt response and late times by nonlinear tails Berti et al. (2025); Leaver (1986); Price (1972); Cheung et al. (2024); Zhu et al. (2024b); De Amicis et al. (2024). In general, to isolate the regime where the QNM model is valid, one needs to apply a window function $W(t)$ to the time-domain signal that removes the early and late time contributions. As a result, the frequency-domain response corresponds to a convolution of Eq. (II.1) with the Fourier transform of the window $W(f)$.
@@
Since the choice of window is not unique, in Fig. 1 we consider a typical ringdown system and compare the frequency-domain waveforms obtained with the standard Heaviside window and with the mirroring technique adopted in our work. Since the Heaviside filter and the mirroring produce a discontinuity at $t=t0$ respectively in the strain and in its first derivative, the corresponding frequency falloffs are $1/f$ and $1/f^{2}$. We find that with the Heaviside window, there is significant spectral leakage from the 220 mode, which contaminates the peaks of the higher harmonics, such as the 330 and 550 modes shown in Fig. 1. With the mirroring, the 330 mode is much less contaminated, due to suppressed spectral leakage from the 220. However, due to its smaller amplitude, the 550 is still significantly contaminated. More in general, depending on the choice of window function, spectral leakage can contaminate the frequency-domain description of the sub-dominant higher frequency modes.
@@ II.3 Modes
{tblr}
colspec=lllllll, row1=font=, column1=font=, roweven=bg=gray!10, cells=c
& Linear Quadratic Fundamental Overtone Prograde Retrograde
220 ✓ ✓ ✓
210 ✓ ✓ ✓
330 ✓ ✓ ✓
320 ✓ ✓ ✓
440 ✓ ✓ ✓
550 ✓ ✓ ✓
[…]
211 ✓ ✓ ✓
331 ✓ ✓ ✓
[…]
r210 ✓ ✓ ✓
[…]
220$\times$ 330 ✓ ✓ ✓
@@
In our work, we use the frequency domain waveform of Eq. (II.1) with the 13 modes listed in Table II.3 as the “true” ringdown signal. These modes can be classified according to the following three criteria:
  • Order in perturbation theory. We consider both linear and quadratic modes. Linear modes can be labeled with the angular and overtone numbers $\ell mn$. We choose to label quadratic modes by $(\ell_{1}m_{1}n_{1})\times(\ell_{2}m_{2}n_{2})$, where $\ell_{i}$ and $m_{i}$ are the angular numbers of the linear modes that source them. We will only consider quadratic modes with angular number given by $\ell=\ell_{1}+\ell_{2}$ and $m=m_{1}+m_{2}$. The frequencies and damping timescales of quadratic modes are given by $\omega_{\ell mn}=\omega_{\ell_{1}m_{1}n_{1}}+\omega_{\ell_{2}m_{2}n_{2}}$ and $\tau_{\ell mn}^{-1}=\tau_{\ell_{1}m_{1}n_{1}}^{-1}+\tau_{\ell_{2}m_{2}n_{2}}^{-1}$.
  • Overtone number. We only consider modes with $n=0$ (fundamental modes) or $n=1$ (overtone modes).
@@
$\begin{split}&\omega_{\ell mn}=f_{1}+f_{2}(1-\chi_{\rm rem})^{f_{3}}\,;\\ &Q_{\ell mn}=q_{1}+q_{2}(1-\chi_{\rm rem})^{q_{3}}\,;\\ &\tau_{\ell mn}=\frac{2Q_{\ell mn}}{\omega_{\ell mn}}\,,\end{split}$ (12)
@@ III Assessing waveform inaccuracy
In the following, for completeness, we review the derivation for the different errors, along with the waveform accuracy criteria. We set the GW data in a given detector to be $d(t)=h_{\rm GR}(\bm{\theta}_{\rm tr};t)+n(t)$, where $n(t)$ is the detector noise. For stationary and Gaussian noise, the likelihood $p(d|\bm{\theta})$ corresponding to the approximate template $h_{\rm AP}$ is then given by
@@
[…]
@@
$\Delta\theta^{i}&=\Delta^{(n)}\theta^{i}+\Delta^{\text{(Sys)}}\theta^{i},\\ \Delta^{(n)}\theta^{i}&=\left(\Gamma^{-1}\right)^{ij}\left(\partial_{j}h_{\text{AP}}\left|\right.n\right)\Big{|}_{\theta^{i}=\theta^{i}_{\rm tr}},\\ \Delta^{\text{(Sys)}}\theta^{i}&=\left(\Gamma^{-1}\right)^{ij}\left(\partial_{j}h_{\text{AP}}\left|\right.h_{\text{GR}}-h_{\text{AP}}\right)\Big{|}_{\theta^{i}=\theta^{i}_{\rm tr}}.\end{aligned}$ (20)
@@
$\begin{split}\langle&\Delta^{(n)}\theta^{i}\Delta^{(n)}\theta^{j}\rangle=\\ &=\langle\left(\Gamma^{-1}\right)^{ik}\left(\Gamma^{-1}\right)^{jl}\left(\partial_{k}h_{\text{AP}}\left|\right.n\right)\left(n\left|\right.\partial_{l}h_{\text{AP}}\right)\rangle\left.\right|_{\theta=\theta_{\text{tr}}}=\\ &=\left(\Gamma^{-1}\right)^{ij}\left.\right|_{\theta=\theta_{\text{tr}}}\,,\end{split}$ (21)
@@
Requiring Eq. (23) for all $\theta^{i}$ translates to Chatziioannou et al. (2017); Chandramouli et al. (2025)
@@
where $1-\mathcal{M}$ is the mismatch, $D$ is the number of parameters of the model $h_{\rm AP}$, and $\rho$ is the optimal SNR. In the mismatch criterion given by Eq. (25), we have neglected the contribution from the fitting factor, which makes the criterion conservative Flanagan and Hughes (1998); Lindblom et al. (2008); Chatziioannou et al. (2017); Pürrer and Haster (2020); Chandramouli et al. (2025); Toubiana and Gair (2024). In other words, when Eq. (25) is satisfied, the approximate waveform is sufficiently accurate.
@@ IV Results
[…]
@@
In the context of the ringdown, we will consider $h_{{\rm GR}}$ to be the template given by Eq. (II.1) including all the 13 modes from the set described in Sec. II.3. On the other hand, we will consider a family of approximate templates $h_{\rm AP}^{(N)}$ including a number $N<13$ of modes. Recall that the set of parameters $\{\theta^{i}\}$ of $h_{\text{GR}}$ includes all the amplitudes and the phases of the $N$ modes, the logarithm of the remnant mass $\log M$, the spin of the remnant BH $\chi_{\rm rem}$, and the four angles $\theta$, $\phi$, $\psi$ and $\iota$, while we restrict ourselves to the case in which $\varphi=0$, amounting to $D=2(N+3)$ free parameters. Notice that we do not include the luminosity distance $d_{L}$ in the Fisher matrix, as it is completely degenerate with the mode amplitudes. Specifically, only the effective amplitudes $A_{\ell mn}^{\text{eff}}=A_{\ell mn}/d_{\ell}$ are measurable. Rather than working with $A_{\ell mn}^{\text{eff}}$, we choose to fix $d_{\ell}$ to its true value, in order to keep the amplitudes $\mathcal{O}(1)$. For the same reason, we do not include the angle $\varphi$, which has no impact on the observables besides shifting the mode phases $\phi_{\ell mn}$. In practice, we set $\varphi=0$ for all systems.
In order to study the dependence of the errors on the number of modes $N$, we adopt the SNR of single-mode template as a criterion for ordering the QNMs. The SNR of each mode, and the resulting hierarchy of the modes, are strongly dependent on the mass ratio and the spin. For instance, highly mass-symmetric systems result in faint $-\ell$ modes, while high spins can affect the spectrum in different ways, e.g. enhancing the contribution from retrograde modes. Additionally, the angles, in particular $\iota$, can influence the luminosity of the QNMs.
@@ V Conclusions
[…]
@@
We would also like to emphasize that, while the use of the low-frequency approximation — compared to the full LISA response — may have a limited effect on the results, the omission of Time Delay Interferometry (TDI) can have a larger impact. This is particularly relevant for low-mass sources, whose ringdown phase may be shorter than the light travel time across the LISA arms. In general, the TDI response overlaps signals from different times, mixing pre-merger, merger and ringdown components. In order to isolate the ringdown phase and mitigate this issue, the analysis could be started at times later than approximately 68 s  [^5] This corresponds to the light travel time for a second-generation TDI configuration. after the luminosity peak. However, this delay can span the entire ringdown duration for some low-mass sources. Therefore, new methods for analyzing ringdown signals are needed to account for complex detector responses, such as that of LISA.

### v2

@@ Abstract
Inaccurate modeling of gravitational-wave signals can introduce systematic biases in the inferred source parameters. As detector sensitivities improve and signals become louder, mitigating such waveform-induced systematics becomes increasingly important. In this work, we assess the systematic biases introduced by an incomplete description of the ringdown signal from massive black hole binaries in the LISA band. Specifically, we investigate the impact of mode truncation in the ringdown template. Using a reference waveform composed of 13 modes, we establish a mode hierarchy and determine the minimum number of modes required to avoid parameter biases across a wide range of LISA sources. For typical systems with masses $\sim 10^{6}$$10^{7}\,M_{\odot}$ at redshifts $z\sim 2$66, we find that at least 3–6 modes are needed for accurate parameter estimation, while high-SNR events may need at least 10 modes. Our results are a window-insensitive lower bound on the minimum number of modes, as more modes may be needed depending on the choice of time-domain windowing of the post-merger signal.
@@ I Introduction
As a binary black hole (BH) system emits gravitational waves (GWs), the two black holes spiral closer together, ultimately merging in a coalescence event. Immediately after the merger, the remnant BH is in a perturbed state and continues to emit GWs in a ringdown phase, gradually settling into a stable configuration. The GW signal in this regime is well understood in terms of BH perturbation theory. In general, any perturbation of the spacetime metric of a rotating BH in General Relativity (GR), at linear order, is governed by the Teukolsky equation Teukolsky (1973). Solutions to the Teukolsky equation, given appropriate boundary conditions, are expressed as a linear superposition of quasi-normal modes (QNMs), i.e. exponentially damped sinusoids Vishveshwara (1970); Chandrasekhar and Detweiler (1975); Anninos et al. (1993); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). Specifically, QNMs are characterized by complex frequencies and amplitudes, with the imaginary part of the frequency determining the damping timescale.
@@ II.1 Time and frequency-domain models
where $M$ is the remnant BH mass in the detector frame, $d_{L}$ is the luminosity distance to the source, $\omega_{\ell mn}$ and $1/\tau_{\ell mn}$ are the real and imaginary parts of the QNM frequency, and $(\iota,\varphi)$ are the angular coordinates of the direction of propagation as seen from the remnant BH, with $\prescript{}{-2}{Y_{\ell m}}$ the spin-weighted spherical harmonics. Note that for a given set of $(\ell mn)$, there are two mode contributions $\mathcal{A}_{\ell mn}$ and $\mathcal{A}^{\prime}_{\ell mn}$: this is because for any given QNM with frequency $\omega_{\ell mn}$ and damping time $\tau_{\ell mn}$, there exists a “mirror mode” also solving the equations of motion, which corresponds to the second term in Eq. (II.1Berti et al. (2009, 2006b). Finally, for this study, we set the starting time to $t_{0}=20M$ after the luminosity peak. For further details on the choice of the starting time, see App. B
@@
where * denotes complex conjugation. This property follows from equatorial symmetry Bucciotti et al. (2024a); Isi and Farr (2021), i.e. from neglecting orbital precession in the merger waveform. [^1] This may be only approximately true for precessing sources, but a full characterization of the amplitude of the mirror modes for precessing systems has not yet been considered in the literature.
@@
$\times\sum_{\ell mn}\frac{A_{\ell mn}Y_{\ell m}^{+,\times}}{\tau_{\ell mn}}\left(L_{\ell mn}^{+}\pm L_{\ell mn}^{-}\right)\,$, (4)
@@
Assuming that the detector response is time-independent, and given two times series $u(t)$ and $v(t)$, we define the matched filter inner product as
@@ II.2 Validity regime of the QNM model
To assess the number of modes needed to ensure waveform accuracy, it is important to account for the validity regime of the frequency-domain QNM model given by Eq. (II.1). The time-domain damped-sinusoid QNM model given by Eq. (II.1) starts being valid only a short time after the peak of the strain amplitude, because early times are dominated by the prompt response, and ceases to be valid at late times, where nonlinear tails become significant Berti et al. (2025); Leaver (1986); Price (1972); Cheung et al. (2024); Zhu et al. (2024b); De Amicis et al. (2024). In general, to isolate the regime where the QNM model is valid, one needs to apply a window function $W(t)$ to the time-domain signal that removes the early and late time contributions. As a result, the frequency-domain response corresponds to a convolution of Eq. (II.1) with the Fourier transform of the window $W(f)$.
@@
Since the choice of window is not unique, in Fig. 1 we consider a typical ringdown system and compare the frequency-domain waveforms obtained with the standard Heaviside window and with the mirroring technique adopted in our work. Since the Heaviside filter and the mirroring produce a discontinuity at $t=t_{0}$ respectively in the strain and in its first derivative, the corresponding frequency falloffs are $1/f$ and $1/f^{2}$. We find that with the Heaviside window, there is significant spectral leakage from the 220 mode, which contaminates the peaks of the higher harmonics, such as the 330 and 550 modes shown in Fig. 1. With the mirroring, the 330 mode is much less contaminated, due to suppressed spectral leakage from the 220. However, due to its smaller amplitude, the 550 is still significantly contaminated. More in general, depending on the choice of window function, spectral leakage can contaminate the frequency-domain description of the sub-dominant higher frequency modes.
@@ II.3 Modes
Linear QuadraticQuadratic Fundamental Overtone Prograde Retrograde
220
210
330
320
440
550
221
211
331
r220
r210
$\times 220$ $\checkmark$
@@
In our work, we use the frequency domain waveform of Eq. (II.1) with the 13 modes listed in Table 1 as the “true” ringdown signal. These modes can be classified according to the following three criteria:
  • Order in perturbation theory. We consider both linear and quadratic modes. In fact, quadratic QNMs have recently been shown to be potentially detectable with LISA Yi et al. (2024). Linear modes can be labeled with the angular and overtone numbers $\ell mn$. We choose to label quadratic modes by $(\ell_{1}m_{1}n_{1})\times(\ell_{2}m_{2}n_{2})$, where $\ell_{i}$ and $m_{i}$ are the angular numbers of the linear modes that source them. We will only consider quadratic modes with angular number given by $\ell=\ell_{1}+\ell_{2}$ and $m=m_{1}+m_{2}$. The frequencies and damping timescales of quadratic modes are given by $\omega_{\ell mn}=\omega_{\ell_{1}m_{1}n_{1}}+\omega_{\ell_{2}m_{2}n_{2}}$ and $\tau_{\ell mn}^{-1}=\tau_{\ell_{1}m_{1}n_{1}}^{-1}+\tau_{\ell_{2}m_{2}n_{2}}^{-1}$.
  • Overtone number. We only consider modes with $n=0$ (fundamental modes) or $n=1$ (overtone modes). Higher overtone modes are typically too short-lived to be relevant for our analysis.
@@
[…]
@@ III Assessing waveform inaccuracy
In the following, for completeness, we review the derivation for the different errors, along with the waveform accuracy criteria. We set the GW data in a given detector to be $d(t)=h_{\rm GR}(\bm{\theta}_{\rm tr};t)+n(t)$, where $n(t)$ is the detector noise. For stationary and Gaussian noise, the likelihood $p(d|\bm{\theta})$ corresponding to the approximate template $h_{\rm AP}$ is then given, up to a constant, by
@@
where $\delta_{D}$ is the Dirac delta function and $S_{n}$ is the PSD. The symbol $\langle\,\cdot\,\rangle$ indicates the ensemble average over all possible noise realizations. For Gaussian and stationary noise, note that Eq. (14) also implies the completeness relation given by
@@
$\begin{aligned} \Delta\theta^{i}&=\Delta^{(n)}\theta^{i}+\Delta^{\text{(Sys)}}\theta^{i},\\ \Delta^{(n)}\theta^{i}&=\left(\Gamma^{-1}\right)^{ij}\left(\partial_{j}h_{\text{AP}}\left|\right.n\right)\Big|_{\theta^{i}=\theta^{i}_{\rm tr}},\\ \Delta^{\text{(Sys)}}\theta^{i}&=\left(\Gamma^{-1}\right)^{ij}\left(\partial_{j}h_{\text{AP}}\left|\right.h_{\text{GR}}-h_{\text{AP}}\right)\Big|_{\theta^{i}=\theta^{i}_{\rm tr}}.\end{aligned}$ (20)
@@
[…]
@@
Requiring that Eq. (23) holds for all $\theta^{i}$ translates to Chatziioannou et al. (2017); Chandramouli et al. (2025)
@@
where 11-$\mathcal{M}$ is the mismatch, $D$ is the number of parameters of the model $h_{\rm AP}$, and $\rho$ is the optimal SNR. In the mismatch criterion given by Eq. (25), we have neglected the contribution from the fitting factor, which makes the criterion conservative Flanagan and Hughes (1998); Lindblom et al. (2008); Chatziioannou et al. (2017); Pürrer and Haster (2020); Chandramouli et al. (2025); Toubiana and Gair (2024). In other words, when Eq. (25) is satisfied, the approximate waveform is sufficiently accurate.
@@ IV Results
Figure 2: QNMs ordered by SNR from the loudest (up) to the quietest (down). We take 3 representative values of spin and mass ratio, fixing the primary mass to $10^{6}\,M_{\odot}$, the luminosity distance to $5\text{Gpc}$, and the angles to $\theta=\psi=\iota=\pi/3$, $\phi=0$. The odd-$m$ modes are highlighted in blue. Note their absence in a mass-symmetric system, and their increasing relevance as we take a small $q$. Further observe that they tend to climb the ranking as we consider more spinning systems. Besides these two features, for highly spinning and mass-asymmetric systems, overtones tend to become more important. The most peculiar case comes with the $331$, which becomes louder than the fundamental $330$.
@@
In the context of the ringdown, we will consider $h_{{\rm GR}}$ to be the template given by Eq. (II.1) including all the 13 modes from the set described in Sec. II.3. On the other hand, we will consider a family of approximate templates $h_{\rm AP}^{(N)}$ including a number $N<13$ of modes. Recall that the set of parameters $\{\theta^{i}\}$ of $h_{\text{GR}}$ includes all the amplitudes and the phases of the $N$ modes, the logarithm of the remnant mass $\log M$, the spin of the remnant BH $\chi_{\rm rem}$, and the four angles $\theta$, $\phi$, $\psi$ and $\iota$, while we restrict ourselves to the case in which $\varphi=0$, amounting to $D=2(N+3)$ free parameters. Notice that we do not include the luminosity distance $d_{L}$ in the Fisher matrix, as it is completely degenerate with the mode amplitudes. Specifically, only the effective amplitudes $A_{\ell mn}^{\text{eff}}=A_{\ell mn}/d_{L}$ are measurable. Rather than working with $A_{\ell mn}^{\text{eff}}$, we choose to fix $d_{L}$ to its true value, in order to keep the amplitudes $\mathcal{O}(1)$. For the same reason, we do not include the angle $\varphi$, which has no impact on the observables besides shifting the mode phases $\phi_{\ell mn}$. In practice, we set $\varphi=0$ for all systems.
In order to study the dependence of the errors on the number of modes $N$, we adopt the SNR of single-mode template as a criterion for ordering the QNMs. The SNR of each mode, and the resulting hierarchy of the modes, are strongly dependent on the mass ratio and the spin. For instance, highly mass-symmetric systems result in faint odd-$\ell$ modes, while high spins can affect the spectrum in different ways, e.g. enhancing the contribution from retrograde modes. Additionally, the angles, in particular $\iota$, can influence the luminosity of the QNMs.
@@
We finally stress that the FIM approach relies on the assumption of high SNR. This requirement is fulfilled by typical massive black hole binaries in the LISA band, which span a wide range of the parameter space shown in Fig. 4. Sources with SNR $\in(8,15)$ correspond to the boundaries of Fig. 4. Our FIM approach thus is not reliable for this (small) region of parameter space with low SNR Vallisneri (2008); Nicholson and Vecchio (1998); Rodriguez et al. (2013); Cokelaer (2008). For these sources, a full Bayesian analysis would be needed to validate the estimation of $N_{\min}$ against our FIM-based prediction.
@@ V Conclusions
Our main result, represented in the waterfall plot of Fig. 4, is the minimum number of modes $N_{\min}$ required for unbiased parameter estimation. As discussed in Sec. III, due to the high-frequency cutoff introduced in our frequency-domain model to mitigate spectral leakage from the dominant mode, the estimate for $N_{\text{min}}$ should be interpreted as a lower bound that is insensitive to the choice of time window, used to isolate the ringdown stage. We determined $N_{\min}$ based on two equivalent waveform accuracy criteria: a direct comparison of systematic and statistical errors within the linear-signal approximation, and, separately, via the mismatch criterion given by Eq. (25) (with both methods in excellent agreement). Our findings highlight the importance of carefully selecting the waveform template across the LISA parameter space. For sources at redshifts around $z\sim 2$66, corresponding to the peak in the merger rate of massive BH binary formation and coalescence models Colpi et al. (2024), we find that a minimum of 3366 modes are needed to avoid systematic biases. This result holds if the modes are included in the template following a single-mode SNR ranking, starting from the loudest for a given system. Note that the presence of each mode depends on the excitation amplitude in the pre-ringdown phase, therefore, the observation of the QNMs is highly linked to the configuration of the system. For massive BH binary sources at low redshifts of $z<1$, as many as $10$ modes may be required to ensure unbiased parameter estimation.
@@
We would also like to emphasize that, while the use of the low-frequency approximation — compared to the full LISA response — may have a limited effect on the general results, the omission of Time Delay Interferometry (TDI) can have a larger impact. This is particularly relevant for low-mass sources, whose ringdown phase may be shorter than the light travel time across the LISA arms. In general, the TDI response overlaps signals from different times, mixing pre-merger, merger and ringdown components. In order to isolate the ringdown phase and mitigate this issue, the analysis could be started at times later than approximately 68 s  [^6] This corresponds to the light travel time for a second-generation TDI configuration. after the luminosity peak. However, this delay can span the entire ringdown duration for some low-mass sources. Therefore, new methods for analyzing ringdown signals are needed to account for complex detector responses, such as that of LISA.